# Linking discrete and continuous models of cell birth and migration

### v2

@@ 2.1 Models of migration
Here, $\textbf{F}^{(i)}_{\text{int}}$ is the net force arising from all cell–cell interactions according to the potential $W_{\text{MM}}(\textbf{r})$, where r is the inter-particle displacement. The potential can encode cell–cell repulsion and adhesion, depending on the sign of its gradient along the direction between two cell centres. Many choices for $W_{\text{MM}}(\textbf{r})$ are possible, including harmonic, power-law, Morse, and Leonard-Jones potentials, among others. Here we use exponential potentials given by:
@@ 2.4 Parameter estimation procedure
We consider biologically meaningful time scales (i.e., days), length scales (i.e., mm), and cell densities and stress empirical units throughout our results. This choice supports future studies that may treat pattern formation with multiple cell types. Throughout our simulations, we consider a domain of size $3$ mm $\times$ 3 mm (with one $1$D exception in Fig. 7). We implement four initial conditions to extract common features of cell interactions from different geometric scenarios: a square region of melanophores in the centre of the domain (“Box”), which mimics the symmetry of the domain; a single stripe of melanophores (“Stripe”), which is motivated by the typical patterning observed in wild-type zebrafish; two rectangular regions of melanophores (“Offset rectangles”), which take into account non-standard geometries and the meeting of two disjoint melanophore populations; and two melanophore stripes (“Two stripes”), which explores the interactions between two disjoint melanophore populations with biologically realistic sharp fronts, see Supplementary Fig. 6 for a summary of these initial conditions. We initialise individual ABM simulations by sampling cell positions uniformly in these regions for each respective initial condition, and initialise our continuous models by setting the cell density uniformly equal to the estimated biological density of $400$ cells/mm2 [21].
@@ 3.3 Cell movement and proliferation
Fig. 10(j), which depicts the time evolution of the estimated radius of support for the PDE and EA ABM results, shows that a reasonably accurate continuous description of the combined model can be obtained by using the scaling parameters obtained from a modular approach. The supports of the ABM and PDE solutions both increase at roughly the same rate, although the degree to which the solutions agree can be affected by $N_{\text{bir}}$: when this parameter is small, the PDE solution travels at a faster rate than the ABM solution, whereas the opposite occurs when this value is large (greater than $N_{\text{bir}}=50$ positions/day). At intermediate values of $N_{\text{bir}}$ (i.e., $N_{\text{bir}}=50$ positions/day), however, the ABM and PDE solution curves are almost identical. Fig. 10(i), which presents the number of cells over time, yields similar observations: the ABM and PDE solutions exhibit similar dynamics over the time period investigated here and there are certain values of $N_{\text{bir}}$ for which the solution curves are nearly identical. Fig. 11 further demonstrates that these observations do not depend on the specific choice of initial condition. Fig.s 10 and 11 demonstrate that combining movement with proliferation also dissipates the oscillatory bands that we observed for movement alone in Fig. 5. This is likely because the stochastic addition of cells in the birth model disrupts the regular cell spacing created by the movement model. Furthermore, the EA ABM and PDE solutions exhibit similar characteristic profiles without regions of high cell density around the edge of the initial condition support, in contrast to the birth-only model (Fig. 8).
@@ 4 Discussion
Our model and parameter-fitting code is publicly available on GitHub [94]. Supplementary figures and videos are available on Figshare at https://figshare.com/projects/Linking_discrete_and_continuous_models_of_cell_birth_and_migration_in_one_population/171234. Data used to generate figures are given in Dryad https://datadryad.org/stash/share/ErRoj5orc7qjK2NuuKSp3rWJvNz6QPLSNhPlmlO52BU.

### v3

@@ 2.1 Models of migration
Here, $\textbf{F}^{(i)}_{\text{int}}$ is the net force arising from all cell–cell interactions according to the potential $W_{\text{MM}}(\textbf{r})$, where r is the inter-particle displacement. The potential can encode cell–cell repulsion and adhesion, depending on the sign of its gradient along the direction between two cell centres. Many choices for $W_{\text{MM}}(\textbf{r})$ are possible, including harmonic, power-law, Morse, and Lennard-Jones potentials, among others. Here we use exponential potentials given by:
@@ 2.4 Parameter estimation procedure
We consider biologically meaningful time scales (i.e., days), length scales (i.e., mm), and cell densities and stress empirical units throughout our results. This choice supports future studies that may treat pattern formation with multiple cell types. Throughout our simulations, we consider a domain of size $3$ mm $\times$ 3 mm (with one $1$D exception in Fig. 7). We implement four initial conditions to extract common features of cell interactions from different geometric scenarios. The first involves a square region of melanophores in the centre of the domain (“Box”), which mimics the symmetry of the domain. For the second initial condition, we place a single stripe of melanophores (“Stripe”), which is motivated by the typical patterning observed in wild-type zebrafish. For the third initial condition, we consider two rectangular regions of melanophores (“Offset rectangles”), which take into account non-standard geometries and the meeting of two disjoint melanophore populations. Finally, the last initial condition we consider involves two melanophore stripes (“Two stripes”), which explores the interactions between two disjoint melanophore populations with biologically realistic sharp fronts. (See Supplementary Fig. 6 for a summary of these initial conditions.) We initialise individual ABM simulations by sampling cell positions uniformly in these regions for each respective initial condition, and initialise our continuous models by setting the cell density uniformly equal to the estimated biological density of $400$ cells/mm2 [21].
@@ 3.3 Cell movement and proliferation
Fig. 10(j), which depicts the time evolution of the estimated radius of support for the PDE and EA ABM results, shows that a reasonably accurate continuous description of the combined model can be obtained by using the scaling parameters obtained from a modular approach. The supports of the ABM and PDE solutions both increase at roughly the same rate, although the degree to which the solutions agree can be affected by $N_{\text{bir}}$: when this parameter is small, the PDE solution travels at a faster rate than the ABM solution, whereas the opposite occurs when this value is large (greater than $N_{\text{bir}}=50$ positions/day). At intermediate values of $N_{\text{bir}}$ (i.e., $N_{\text{bir}}=50$ positions/day), however, the ABM and PDE solution curves are almost identical. Fig. 10(i), which presents the number of cells over time, yields similar observations: the ABM and PDE solutions exhibit similar dynamics over the time period investigated here and there are certain values of $N_{\text{bir}}$ for which the solution curves are nearly identical. Fig. 11 further demonstrates that these observations do not depend on the specific choice of initial condition. Figs. 10 and 11 demonstrate that combining movement with proliferation also dissipates the oscillatory bands that we observed for movement alone in Fig. 5. This is likely because the stochastic addition of cells in the birth model disrupts the regular cell spacing created by the movement model. Furthermore, the EA ABM and PDE solutions exhibit similar characteristic profiles without regions of high cell density around the edge of the initial condition support, in contrast to the birth-only model (Fig. 8).
@@ 4 Discussion
Our model and parameter-fitting code is publicly available on GitHub [94]. Supplementary figures and videos are available on Figshare at https://figshare.com/projects/Linking_discrete_and_continuous_models_of_cell_birth_and_migration_in_one_population/171234. Data used to generate figures are given in Dryad https://doi.org/10.5061/dryad.s4mw6m9cb.